An college savings fund will be worth $34,000 in 10 years. Find the present value if interest is at 4.4% compounded continuou3. (7 points) Find the area of the region bounded by the graphs of y = 3x2 + 2x + 2, y = -2x - 3, x= -1 and x = 4.
Answer #1

Date : -22+2 - 2x-3. 12=-1 and 2=4. 23 +21173 350 224 Anea a Shaded region Y=-22-3 A = (y of paraboola- y of line) de (3x²+2Р- зЧ, ооо , с. 4. 1. – 104 сом 34, оор – P. + чу 100 31 000 - Р. +0. оче 6 – 34 доое (1+0.4ч )о зч, боо ( +0. Чч ) | -